# MT-Patcher: Selective and Extendable Knowledge Distillation from Large Language Models for Machine Translation

### v1

@@ Abstract
[…]
*** This work is done while Jiahuan’s internship at Bytedance.$\dagger$ Corresponding author.
@@ 3.1 Knowledge Selection via Feedbacker
When transferring knowledge from LLMs to existing MT models, traditional SeqKD would finetune the student model on all teacher’s output, ignoring the fact that the student model can already translate most of the examples well. Furthermore, several recent studies have unveiled emergent abilities in LLMs, such as Self-Refinement (Madaan et al., 2023) and Self-Debug (Chen et al., 2023), suggesting that iterative refinement of an initial draft may be a more effective strategy to tap into the knowledge reserves of LLMs.
@@ 6 Conclusion
Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang is the corresponding author. This work is supported by National Science Foundation of China (No. 62376116, 62176120), the Liaoning Provincial Research Foundation for Basic Research (No. 2022-KF-26-02).
@@ Appendix E Prompts for Evaluation
Assuming you are a highly proficient translator skilled at providing detailed and comprehensive assessments of machine translations. I will give you a <srclang> sentence X and its <tgtlang> translation Y, and I would like you to help assess the translation. 1. You should first provide an overall assessment. 2. Following that, - If there are no errors, just say "No error." and do not provide an explanation. - If there are errors, please specify - the error type, - the corresponding segment in the <srclang> sentence X, - the corresponding segment in the translation Y, - the reason for the error, - and the correct translation for the segment - If there are errors, you should also provide a good translation at the end of the assessment. 4. For multiple errors, you should address them separately. 5. Try to pinpoint the smallest segments containing errors and explain them, avoiding cases where the error encompasses the entire sentence. 6. Carefully read the original text and the translation to identify all translation errors. 7. Your response should be in English. 8. Be concise. Now, please assess the following translation: <srclang>: <srctext> <tgtlang>: <tgttext> Assessment:
@@
Suppose you are a language expert of <srclang> and <tgtlang>. Given a sentence X, please point out its topic, domain and style. Input: X: <srctext> Output:
@@
Suppose you are a language expert of <srclang> and <tgtlang>. Given a topic, a domain and a style, as well as a bilingual word pair, please generate a pair of parallel sentences that adhere to the given topic, domain and style. They should also contain the given word pair. Input: Domain: <domain> Topic: <topic> Style: <style> Word Pair: <wordpair> Output:
@@
Assume you are a <srclang> and <tgtlang> language expert with a wealth of knowledge and associative ability in both languages. I will give you a word/phrase P from an <srclang> sentence X. Please associate from the following aspects and generate three words similar to X for each aspect, and provide the <tgtlang> translation of these words. Aspects of association: - Category. What kind of category does this word belong to? - Semantics. What words often appear in the same context as the given word? NOTE, the associated words should be rare words, so that it is unlike for a machine translation system to translate it correctly. Input: X: <srctext> P: <errorword> Output:
@@
Assume you are a language expert in English and Chinese. I will give you a Chinese idiom S, a sentence X that contains S, and a machine-generated English translation Y of the source sentence X. I will also give you the explanation/definition E of the idiom S. Your task is to first identify the translation of S in Y, and judge whether the translation of the idiom is correct. Note: 1. The score range is 0/1/2/3/4/5, where - 0: Completely incorrect translation or no translation - 1: Literal translation of the original, without conveying any implied meaning, leaving non-Chinese background readers baffled - 2: Literal translation of the original, partially conveying the implied meaning, easy for non-Chinese background readers to understand - 3: Interpretative translation of the idiom, but only partially conveying the implied meaning - 4: Interpretative translation of the idiom, fully conveying the implied meaning - 5: The translation perfectly conveys the implied meaning of the idiom, is very easy for all readers to understand, and also maintains the aesthetic sense of the original 2. You should generate the explanation of your decision concisely. Now, please process the following inputs:
@@
Assume you are a language expert in Chinese and English. I will give you a sentence X, the word P in that sentence, and two translations of the sentence X: A and B. Your task is to assess which translation contains the correct translation of the word P. Requirements: (1) Ignore other differences between the two translations. Only compare the translation of the word P. (2) Your answer should first state the reason for your comparison, and then give your comparison. (3) Your comparison should be A, B, C and D. - A: the first translation of the word P is better. - B: the second translation of the word P is better. - C: Both are fine. - D: Both are bad. Now, please process the following inputs:

### v2

@@ Abstract
Jiahuan Li♣∗, Shanbo ${}^{\spadesuit}\dagger$, Shujian ${}^{\clubsuit}\dagger$and Jiajun Chen National Key Laboratory for Novel Software Technology, Nanjing University, China ByteDance Research lijh@smail.nju.edu.cn, {huangsj,chenjj}@nju.edu.cn chengshanbo@bytedance.com
* Work done while Jiahuan Li’s internship at ByteDance Research.$\dagger$ Corresponding authors.
@@ 3.1 Knowledge Selection via Feedbacker
When transferring knowledge from LLMs to existing MT models, traditional SeqKD would finetune the student model on all teacher’s output, ignoring the fact that the student model can already translate most of the examples well. Furthermore, several recent studies have unveiled emergent abilities in LLMs, such as Self-Refinement (Madaan et al., 2023) and Self-Debug (Chen et al., 2024), suggesting that iterative refinement of an initial draft may be a more effective strategy to tap into the knowledge reserves of LLMs.
@@ 6 Conclusion
Acknowledgement
We would like to thank the anonymous reviewers for their insightful comments. Shujian Huang and Shanbo Cheng are the corresponding authors. This work is supported by National Science Foundation of China (No. 62376116, 62176120), the Liaoning Provincial Research Foundation for Basic Research (No. 2022-KF-26-02).
@@ Appendix E Prompts for Evaluation
{mdframed}
Assuming you are a highly proficient translator skilled at providing detailed and comprehensive assessments of machine translations. I will give you a <srclang> sentence X and its <tgtlang> translation Y, and I would like you to help assess the translation.
1. You should first provide an overall assessment.
2. Following that,
- If there are no errors, just say "No error." and do not provide an explanation.
- If there are errors, please specify
- the error type,
- the corresponding segment in the <srclang> sentence X,
- the corresponding segment in the translation Y,
- the reason for the error,
- and the correct translation for the segment
- If there are errors, you should also provide a good translation at the end of the assessment.
4. For multiple errors, you should address them separately.
5. Try to pinpoint the smallest segments containing errors and explain them, avoiding cases where the error encompasses the entire sentence.
6. Carefully read the original text and the translation to identify all translation errors.
7. Your response should be in English.
8. Be concise.
Now, please assess the following translation:
<srclang>: <srctext>
<tgtlang>: <tgttext>
Assessment:
@@ 2. Following that,
{mdframed}
Suppose you are a language expert of <srclang> and <tgtlang>. Given a sentence X, please point out its topic, domain and style.
Input:
X: <srctext>
Output:
@@
{mdframed}
Suppose you are a language expert of <srclang> and <tgtlang>. Given a topic, a domain and a style, as well as a bilingual word pair, please generate a pair of parallel sentences that adhere to the given topic, domain and style. They should also contain the given word pair.
Input:
Domain: <domain>
Topic: <topic>
Style: <style>
Word Pair: <wordpair>
Output:
@@
{mdframed}
Assume you are a <srclang> and <tgtlang> language expert with a wealth of knowledge and associative ability in both languages. I will give you a word/phrase P from an <srclang> sentence X. Please associate from the following aspects and generate three words similar to X for each aspect, and provide the <tgtlang> translation of these words.
Aspects of association:
- Category. What kind of category does this word belong to?
- Semantics. What words often appear in the same context as the given word?
NOTE, the associated words should be rare words, so that it is unlike for a machine translation system to translate it correctly.
Input:
X: <srctext>
P: <errorword>
Output:
@@
{mdframed}
Assume you are a language expert in English and Chinese. I will give you a Chinese idiom S, a sentence X that contains S, and a machine-generated English translation Y of the source sentence X. I will also give you the explanation/definition E of the idiom S. Your task is to first identify the translation of S in Y, and judge whether the translation of the idiom is correct.
Note:
1. The score range is 0/1/2/3/4/5, where
- 0: Completely incorrect translation or no translation
- 1: Literal translation of the original, without conveying any implied meaning, leaving non-Chinese background readers baffled
- 2: Literal translation of the original, partially conveying the implied meaning, easy for non-Chinese background readers to understand
- 3: Interpretative translation of the idiom, but only partially conveying the implied meaning
- 4: Interpretative translation of the idiom, fully conveying the implied meaning
- 5: The translation perfectly conveys the implied meaning of the idiom, is very easy for all readers to understand, and also maintains the aesthetic sense of the original
2. You should generate the explanation of your decision concisely.
Now, please process the following inputs:
@@ 1. The score range is 0/1/2/3/4/5, where
{mdframed}
Assume you are a language expert in Chinese and English. I will give you a sentence X, the word P in that sentence, and two translations of the sentence X: A and B. Your task is to assess which translation contains the correct translation of the word P.
Requirements:
(1) Ignore other differences between the two translations. Only compare the translation of the word P.
(2) Your answer should first state the reason for your comparison, and then give your comparison.
(3) Your comparison should be A, B, C and D.
- A: the first translation of the word P is better.
- B: the second translation of the word P is better.
- C: Both are fine.
- D: Both are bad.
Now, please process the following inputs: